OpenStudy (anonymous):

Solve, using the substitution method. 2x + 4y = 16 x – 7 = – 4y A.(0,-7) B.(0, 7/4) C.(9, -1/2) D.(17, -5/2)

OpenStudy (danjs):

x = -4y + 7 put into first equation

OpenStudy (danjs):

2(-4y+7) +4y = 16

OpenStudy (danjs):

-8y + 14 + 4y = 16 -4y = 2 y = -1/2

OpenStudy (anonymous):

x=9?

OpenStudy (danjs):

Put that y value into one of the equations in the question.

OpenStudy (danjs):

x - 7 = -4(-1/2) = 2 x=9 yep

OpenStudy (anonymous):

the answer is C

OpenStudy (danjs):

ok, that was figured out 2hrs ago above.